Suppose f : N → N is a function defined by ƒ(x) = a(x + b), where a, b E N. If f is a bijective function,




then find the value of a + b.

1
Expert's answer
2022-02-15T10:33:45-0500

Solution:

Since the function is bijective, we know that it is injective and surjective.

Thus define a function;

"\displaystyle
f:\N\rightarrow\N,\text{ defined by }f(x)=\begin{cases}x-1 & \text{if }x\text{ is even}\\
x+1 & \text{if }x\text{ is odd}\end{cases}"

The above function is bijective.

Thus, we have "\displaystyle
a=1\in\N" in both cases and "\displaystyle
b=\begin{cases}-1 & \text{if }x \text{ is even}\\1\in\N&\text{if }x \text{ is odd}\end{cases}"

Hence, for this function with "\displaystyle
a,b\in \N\text{ we have }a+b=2"